OpenStudy (anonymous):

in the equation: -4x + 12 = -11 what is the value x?

OpenStudy (anonymous):

Remove the +12 from the equation. Do so by subtracting it from both sides. -4x + 12 - 12 = -11 - 12 -4x = -11 - 12 After you subtract 12 from -11, divide that new number by four, leaving you with x = your answer.

OpenStudy (anonymous):

so x = 5.75?

OpenStudy (anonymous):

What you're trying to do is isolate x. You do this by taking out all other numbers on the side of the equation that x is on. So, if you want to take out positive 12, subtract it; BUT, what you do on one side, you have to do to the other side.

OpenStudy (anonymous):

yea you're right. good job :)

OpenStudy (anonymous):

oh okay so : -4x = -23

OpenStudy (anonymous):

Indubitably.

OpenStudy (anonymous):

okay thanks i just wanted to make sure

OpenStudy (anonymous):

yea, but then simplify by dividing by 4, leaving you just with "x"

OpenStudy (anonymous):

and u also dived the other side by -4 an u get 5.75. x= 5.75

OpenStudy (anonymous):

Thanks ! Bye!